OpenStudy (anonymous):

wha is the vaule of x+99=287

OpenStudy (zepp):

What can you do to isolate x?

OpenStudy (anonymous):

- the 99 from both sides

OpenStudy (zepp):

Yes, then what do you get?

OpenStudy (radar):

Or add -99 to both sides of equation getting: x+99-99=287-99

OpenStudy (anonymous):

nooo. x=287-99 x=188

OpenStudy (zepp):

Lui_Rivera, you didn't take off the 99 the left side ..^ nickymarden has beaten me to it..

OpenStudy (anonymous):

heyy nicky your pretty

OpenStudy (anonymous):

you guys just made it harder for him. You just 'pass' the number to the other side of the equal sign. Just remember to change the sign of the number.

OpenStudy (anonymous):

haha, thank u. lol

OpenStudy (anonymous):

r u single

OpenStudy (anonymous):

nope. lol